HETAV PATEL

Grade 11,

The half-life of a first order reaction is 100 seconds . What is the time required for 90% completion of the reaction?

The half-life of a first order reaction is 100 seconds . What is the time required for 90% completion of the reaction?

Grade:8

1 Answers

Sudhir Bishnoi
6 years ago
For first order reaction t1/2=0.693/K
 where, t1/2 is half life and K is rate constant
 Given half life=100 s, therefore K=0.00693
 
Kt=2.303log10A0/A 
Given A=10/100A0=0.1A0
putting these values in this equation
0.00693×t=2.303log1010
0.00693×t=2.303×1
t=2.303/0.00693
t=332.32 sec apox                    and